OpenStudy (anonymous):

Explain how to find the sum positive five plus negative nineusing chips, a number line, or the rules for adding integers. Give the answer with your explanation.

OpenStudy (anonymous):

Let's use the number line! Start at zero, go to the right 5 units (positive five), then go to the left 9 units (negative nine). Stop and look at the number line, you are at -4 position on the number line. That means the sum of +5 and -9 is -4 5+(−9)=5−9=?

OpenStudy (anonymous):

@caprisha Do you understand?

OpenStudy (anonymous):

yes i got it now i got it right.

OpenStudy (anonymous):

Great! Hope I helped!

OpenStudy (anonymous):

yes thanks hopeful i get the next one